### v2

@@ 3.2 Descriptive statistics
Figure 3.2: Highest values among weekly positive response rates (left) and available sample sizes (right) across 94 counties in the catchment area.
@@
Figure 3.3: Comparisons of demographic distributions between the input data of hospital patients and the target population approximated by linked American Community Survey data in the catchment area.
@@
Figure 3.3 compares the sex, race, and age distributions between the hospital patients ($n=128,\!222$) and the population residing in the catchment area ($N=22,\!320,\!702$). The hospital patients have larger proportions of female, White, and older people than the population, and this sample discrepancy will be adjusted by the poststratification step in MRP.
@@
Figure 3.4: Distributions of geographic characteristics based on the linked American Community Survey data in the catchment area.
Figure 3.4 presents the distributions of geographic characteristics. The catchment area of the hospital patients’ residence covers 848 ZIP codes and has a broad and diverse representation in terms of urban/rural areas, area deprivation status, higher education attainment, employment rate, income, and poverty. We expect that socioeconomic measures at the ZIP level would affect individual behaviors and health and be related to viral transmission. The geographic characteristics would explain the spatial variation. The poststratification uses the population counts by ZIP code but does not adjust the geographic characteristics.
3.3 Model fitting
The interface allows users to specify and fit different models with various choices of individual/geographic covariates and fixed/varying effects. The model fitting is via Bayesian computation with Markov chain Monte Carlo algorithms in Stan. Users can specify prior distributions or choose the default prior setups that are weakly informative. The model outputs include summaries and convergence assessments of the posterior sample of model parameters. We give an example of output from model (2.3.2) in Section A of the Appendix.
3.4 Model diagnostics
The interface compares different models and presents model diagnostic results. We employ the approximate leave-one-out cross-validation (LOO-CV) implemented in the R package loo (Vehtari et al., 2024) and posterior predictive checking (PPC; Gelman et al., 1996). The LOO-CV assesses the posterior predictive performance of Bayesian models and compares different models on expected log predictive density (elpd) for new data.
We have compared the following three models with different mean structure and variance specifications. Model A includes the fixed effects of sex and geographic predictors and varying effects of age, race, time in weeks, and ZIP code. Model B adds high-order interactions, between race and college attainment status, to Model A. Model C removes the ZIP-varying effects from Model A.
Model A: $\beta_{1}+\beta_{2}{\rm male}_{j}+\alpha_{{\rm a}[j]}^{\rm age}+\alpha_{{\rm r}[j]}^{\rm race}+\alpha_{{\rm t}[j]}^{\rm time}+\alpha_{1}{\rm ADI}_{{\rm s}[j]}+\alpha_{2}{\rm college}_{{\rm s}[j]}+\alpha_{3}{\rm employment}_{{\rm s}[j]}+\alpha_{4}{\rm income}_{{\rm s}[j]}+\alpha_{5}{\rm poverty}_{{\rm s}[j]}+$
$\alpha_{6}{\rm urbanicity}_{{\rm s}[j]}+e_{{\rm s}[j]}$
Model B: $\beta_{1}+\beta_{2}{\rm male}_{j}+\alpha_{{\rm a}[j]}^{\rm age}+\alpha_{{\rm r}[j]}^{\rm race}+\alpha_{{\rm t}[j]}^{\rm time}+\alpha_{1}{\rm ADI}_{{\rm s}[j]}+\alpha_{2}{\rm college}_{{\rm s}[j]}+\alpha_{3}{\rm employment}_{{\rm s}[j]}+\alpha_{4}{\rm income}_{{\rm s}[j]}+\alpha_{5}{\rm poverty}_{{\rm s}[j]}+$
$\alpha_{6}{\rm urbanicity}_{{\rm s}[j]}+\alpha_{7}{\rm race*urbanicity}_{{\rm s}[j]}+e_{{\rm s}[j]}$
Model C: $\beta_{1}+\beta_{2}{\rm male}_{j}+\alpha_{{\rm a}[j]}^{\rm age}+\alpha_{{\rm r}[j]}^{\rm race}+\alpha_{{\rm t}[j]}^{\rm time}+\alpha_{1}{\rm ADI}_{{\rm s}[j]}+\alpha_{2}{\rm college}_{{\rm s}[j]}+\alpha_{3}{\rm employment}_{{\rm s}[j]}+\alpha_{4}{\rm income}_{{\rm s}[j]}+\alpha_{5}{\rm poverty}_{{\rm s}[j]}+$
$\alpha_{6}{\rm urbanicity}_{{\rm s}[j]}.$
elpd_diff se_diff Model A 0 0 Model B 2.05-2.05- 2.05 0.86 Model C 4.75-4.75- 4.75 3.20
Table 3.1: Model comparisons with the leave-one-out cross-validation. The two columns show the expected log predictive density difference and its standard error, in each case comparing to Model A.
Figure 3.5: Posterior predictive check comparing replicated positive response rates generated from the estimated model to raw rates by week.
3.5 Estimation results
Figure 3.6: Estimated weekly incidence in the community based on the multilevel regression and poststratification (MRP) in comparison with raw values. The shaded areas represent 95% credible intervals.
MRP stabilizes small group estimates and adjusts for the sample discrepancy within each group. Figure 3.7 presents the estimated incidence for White, Black, and other race categories. Whites tend to have lower infection rates than Black and other racial groups, even though most weekly differences are small and not significant. The model does not include time trends varying across racial groups, i.e., without racial moderation effects. The estimated trends are close to paralleling with similar spike and flat periods. Examining racial differences by week, we expect the differences to be small because of weekly small numbers of tests. When we calculate the cumulative incidences through the study period, Whites are less likely to be infected than Black and other racial groups, which is consistent with the literature findings (Magesh et al., 2021). The same observation of trends applies to the sex and age group estimates, given in Figures B.1 and B.2 of the Appendix.
(a) Observed positivity
(b) Estimated incidence
(c) Standard errors
Figure 3.8: Observed and estimated county-level incidence with standard errors during the week of 01/31/2022-02/06/2022 in the catchment area.
We select one week that observes a spike of infection, 01/31/2022-02/06/2022, and present the county-level estimated incidence with standard error (SE) values in Figure 3.8. The collected test records across counties are sparse, where 32 out of 94 counties do not have any tests and 28 counties have only one test during the selected week. The county with the largest number of tests (230) is where the health system is located. The observed positivity values are unreliable, and 52 out of 62 collected values are zero. Among the top five counties that report the largest numbers of tests and any positive cases during the studied week, the demographic distributions of tested patients are generally similar, with an over-representation of female, White, and older people comparing to the ACS data. The MRP estimates are available for all 94 counties based on the predictions with the ACS data. The model fit summaries in the Appendix A show that the estimated coefficient of urbanicity (defined as the percentage of covered census tracts classified as urban, weighted by tract population) is 0.10-0.10- 0.10 with the 95% CI of (0.21,0.02)(-0.21,0.02)( - 0.21 , 0.02 ), which shows that the urban areas tend to have lower infection rates than the non-urban areas. The multilevel model smooths county-level incidence estimates with a range of 0.1%–5.6%, with a median value of 0.6%. The SE values of the 94 county-wise incidence estimates are between 0.001 and 0.037, and the variation generally increases with the estimated incidence.
3.6 Validation and comparison
Figure 3.9: Comparison of COVID-19 tracking trends estimated by applying multilevel regression and poststratification (MRP) to Michigan hospital test data, the Michigan Wastewater monitored COVID-19 surveillance and weekly percentages of Emergency Department (ED) visits diagnosed as COVID-19 in Michigan.
Our surveillance tool leverages routine hospital testing of asymptomatic patients to provide an early indicator of community disease presence, serving a similar function to wastewater monitoring for SARS-CoV-2 in public sewer systems. By tracking trends over time, both tools can detect increases in SARS-CoV-2 prevalence, thereby alerting health agencies to potential surges in cases and an increased clinical burden. We have compared our results to data from the Michigan Wastewater Dashboard for COVID-19 surveillance (Michigan Department of Health & Human Services, 2025) and to weekly percentages of Emergency Department (ED) visits diagnosed as COVID-19 in Michigan, as reported by the CDC COVID Data Tracker (Centers for Disease Control and Prevention, 2025). Figure 3.9 demonstrates that our estimated trends between March 22, 2020, and October 24, 2022, closely align with wastewater-based surveillance for SARS-CoV-2 shed into Michigan’s public sewer systems. Both surveillance methods capture the spikes in November 2020 and January 2022, as well as the downward trend since August 2022. Notably, our surveillance approach can anticipate increases in ED visit numbers reported to the COVID Data Tracker by approximately one to two weeks. This validation supports findings previously reported in Indiana (Covello et al., 2021; Si et al., 2022).
Our results based on a representation adjustment of routine hospital test records serve a synthetic proxy for random sampling. When available, random-sample testing surveys provide valuable benchmarking data and should be leveraged to calibrate other data sources to ensure population representativeness (Irons and Raftery, 2021; Menachemi, 2020). However, increasing nonresponse rates in these surveys necessitate nonresponse bias adjustments (Si et al., 2023, 2024). For example, Yiannoutsos et al. (2021) have applied a similar method to MRP and adjusted for nonresponse bias in a randomized study of COVID-19 testing in Indiana, the response rate of which is 23.6%. Notably, the trends in calibrated new infection numbers reported by Irons and Raftery (2021) are consistent with those seen in the MRP-adjusted hospital test incidence monitoring (Covello et al., 2021) between March 2020 and March 2021 in Indiana, particularly regarding the capture of infection spikes.
4 Discussion
With generalizability as the goal, the MRP method adjusts for selection bias and stabilizes small group estimates. The user-friendly interface enables application of MRP to both cross-sectional and longitudinal data, facilitating subgroup trend analyses over time. Built with the open-source software Stan, R, and Shiny, the interface supports statistical computation, visualization, and is open-source, regularly maintained, and easily installed for local computation. A secure, privacy-focused web-based version is in development. The interface analyzes aggregated data by poststratification cells, which also facilitates disclosure risk control. The open-source interface is easily accessed and regularly maintained, and will continue to evolve and develop over time.
The interface tracks the epidemic and delivers substantive findings in time, which is demonstrated using the Michigan health system data. Previous work used data from a community hospital in Indiana (Covello et al., 2021), and the results show differences between the states. As shown in our county-level estimates in Michigan, geographic variation can be substantial. Expanding to data from more states will further enhance national generalizability. The interface aims to empower users of varying backgrounds to analyze their own localized data effectively.
With respect to our current sampling method for COVID-19 viral tracking, in accordance with accepted American Society of Anesthesiology standards, all preoperative patients in the hospital system are subjected to surgical risk evaluation. Hence, routine hospital testing is already implemented uniformly across the U.S., increasing the operational feasibility of our proposed surveillance system.
Our approach has several key assumptions. First, we assume sample selection is ignorable, conditional on adjusted demographics and geography. The only factor determining the sample inclusion is the selection for elective surgical procedures. Elective surgery patients may differ from the broader community in unmeasured ways, such as healthcare access or overall health status. However, we need high-quality population data on healthcare use and health measures to adjust these potentially confounding factors. Second, we assume a constant ratio of asymptomatic to symptomatic SARS-CoV-2 infections within any demographic and geographic stratum. We use the estimated incidence based on asymptomatic test records to track the infection trend, but not the magnitude. The ratios may change values with new viral variants and immunity levels that are naturally-acquired or vaccine-induced. We have conducted sensitivity analyses by including the time indicators when the new viral was first detected in Michigan in the model and found that the estimated trends are similar, though the incidence rates have slightly shifted. We have also applied the MRP adjustment to viral IgG testing data of the same group of asymptomatic patients and validated the method using verified clinical metrics of viral and symptomatic disease incidence to show the expected biological correlation of these entities with the timing, rate, and magnitude of seroprevalence (Si et al., 2022). Third, the model-based adjustment is subject to model misspecification. We use a Bayesian binomial model, and our subgroup estimates are robust across different outcome models. It is possible that the model fails to capture some data structure. We suggest users conduct thorough model diagnostics, such as the PPC and LOO-CV in our paper, and result validation. Additionally, the interface’s current focus is infection incidence estimation by subgroup, but it is extensible to other epidemiological parameters (e.g., effective reproduction number, infection fatality ratio), which would require integrating individual-level test data with aggregate case and mortality data using hierarchical Bayesian frameworks. Despite these limitations, the flexibility of the MRP approach allows its use with a variety of sampling methods and data sources, supporting broader applicability beyond the specific settings tested so far. Though post-epidemic testing is currently paused, the interface has broad applicability for other disease monitoring and data analyses with population representation.
Originally developed in response to COVID-19, the MRP interface provides a foundation for broader epidemic surveillance and diverse applications in health and social science research. It accommodates time-varying and cross-sectional data, continuous and binary outcomes, and supports subgroup analyses across demographic and geographic domains. Users can specify models, priors, and poststratification data, and analyze probability sample surveys, non-probability samples, and multiple data sources. Future enhancements will address complex spatiotemporal structures, custom prior distributions, and poststratification with incomplete population data.
In summary, the MRP interface offers a reproducible, extendable framework for statistically valid, high-resolution subgroup estimations, positioning it as a valuable tool for ongoing and future public health research efforts.
Acknowledgments
Appendix A Model fitting results
The Stan fit summaries of Model A are displayed in Table A.1.

### v3

@@ 3.2 Descriptive statistics
Figure 3.2: Highest values among weekly positive response rates (left) and available sample sizes (right) across 94 counties in the catchment area, exemplifying the large geographic variation with sparse data.
@@
Figure 3.3: Comparisons of demographic distributions between the input data of hospital patients and the target population approximated by linked American Community Survey data in the catchment area. The hospital patients over-represent female, White, and older people.
@@
Figure 3.3 compares the sex, race, and age distributions between the hospital patients (n=128,222n=128,\!222) and the population residing in the catchment area (N=22,320,702N=22,\!320,\!702). The hospital patients have larger proportions of female, White, and older people than the population, and this sample discrepancy will be adjusted by the poststratification step in MRP.
Figure A.1 in the Appendix presents the distributions of geographic characteristics. The catchment area of the hospital patients’ residence covers 848 ZIP codes and has a broad and diverse representation in terms of urban/rural areas, area deprivation status, higher education attainment, employment rate, income, and poverty. We expect that socioeconomic measures at the ZIP level would affect individual behaviors and health and be related to viral transmission. The geographic characteristics would explain the spatial variation. The poststratification uses the population counts by ZIP code but does not adjust the geographic characteristics.
3.3 Model fitting
We fit the model via Bayesian computation with Markov chain Monte Carlo in Stan and obtain model fit summaries and convergence assessments of the posterior sample of model parameters. We give an example of output from model (2.3.2) in Section B of the Appendix. With the interface, we can also specify and fit different models with various choices of individual/geographic covariates and varying/non-varying effects under different prior distributions.
3.4 Model diagnostics
We compare different models and present model diagnostic results. We employ the approximate leave-one-out cross-validation (LOO-CV) implemented in the R package loo (loo-R) and posterior predictive checking (PPC; ppc:gelman-96). The LOO-CV assesses the posterior predictive performance of Bayesian models and compares different models on expected log predictive density (elpd) for new data.
We have compared the three models with different mean structure and variance specifications, with included predictors listed in Table 3.1. Model A includes coefficients for sex and geographic predictors and varying effects of age, race, time in weeks, and ZIP code. Model B adds high-order interactions, between race and college attainment status, to Model A. Model C removes the ZIP-varying effects from Model A.
[…]
A $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
B $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
C $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 3.1: Predictors included in each model. A checkmark indicates inclusion.
elpd_diff se_diff Model A 0 0 Model B 2.05-2.05 0.86 Model C 4.75-4.75 3.20
Table 3.2: Model comparisons with the leave-one-out cross-validation. The two columns show the expected log predictive density difference and its standard error, in each case comparing to Model A. The negative elpd_diff values show that Model A has the best predictive performance. The se_diff values support whether the improvement of Model A is substantial.
Figure 3.4: Posterior predictive check comparing replicated positive response rates generated from the estimated model to raw rates by week. The replicates are close to the raw data, showing that the model fits the data well.
3.5 Estimation results
Figure 3.5: Estimated weekly incidence in the community based on the multilevel regression and poststratification (MRP) in comparison with raw values. The shaded areas represent 95% credible intervals. The MRP estimates match the sample demographics with those in the population and are generally higher than the raw positivity, mainly because of the test sensitivity.
MRP stabilizes small group estimates and adjusts for the sample discrepancy within each group. Figure 3.6 presents the estimated incidence for White, Black, and other race categories. Whites tend to have lower infection rates than Black and other racial groups, even though most weekly differences are small and not significant. The model does not include time trends varying across racial groups, i.e., without racial moderation effects. The estimated trends are close to paralleling with similar spike and flat periods. Examining racial differences by week, we expect the differences to be small because of weekly small numbers of tests. When we calculate the cumulative incidences through the study period, Whites are less likely to be infected than Black and other racial groups, which is consistent with the literature findings (Magesh_2021). The same observation of trends applies to the sex and age group estimates, given in Figures C.1 and C.2 of the Appendix.
(a) Observed positivity
(b) Estimated incidence
(c) Standard errors
Figure 3.7: Observed and estimated county-level incidence with standard errors during the week of 01/31/2022–02/06/2022 in the catchment area. The model estimates reduce between-county variation, and their standard errors generally increase with the estimated incidence values.
We select one week that observes a spike of infection, 01/31/2022–02/06/2022, and present the county-level estimated incidence with standard error (SE) values in Figure 3.7. The collected test records across counties are sparse, where 32 out of 94 counties do not have any tests and 28 counties have only one test during the selected week. The county with the largest number of tests (230) is where the health system is located. The observed positivity values are unreliable, and 52 out of 62 collected values are zero. Among the top five counties that report the largest numbers of tests and any positive cases during the studied week, the demographic distributions of tested patients are generally similar, with an over-representation of female, White, and older people compared to the ACS data. The MRP estimates are available for all 94 counties based on the predictions with the ACS data. The model fit summaries in the Appendix A show that the estimated coefficient of urbanicity (defined as the percentage of covered census tracts classified as urban, weighted by tract population) is 0.10-0.10 with the 95% interval of (0.21,0.02)(-0.21,0.02), which shows that the urban areas tend to have lower infection rates than the non-urban areas. The multilevel model smooths county-level incidence estimates with a range of 0.1%–5.6%, with a median value of 0.6%. The SE values of the 94 county-wise incidence estimates are between 0.001 and 0.037, and the variation generally increases with the estimated incidence.
3.6 Validation and comparison
Figure 3.8: Comparison of COVID-19 tracking trends estimated by applying multilevel regression and poststratification (MRP) to Michigan hospital test data, the Michigan wastewater monitored COVID-19 surveillance and weekly percentages of Emergency Department (ED) visits diagnosed as COVID-19 in Michigan, showing an alignment of transmission trends and early warning of clinical burdens.
Our surveillance tool leverages routine hospital testing of asymptomatic patients to provide an early indicator of community disease presence, serving a similar function to wastewater monitoring for SARS-CoV-2 in public sewer systems. By tracking trends over time, both tools can detect increases in SARS-CoV-2 prevalence, thereby alerting health agencies to potential surges in cases and an increased clinical burden. We have compared our results to data from the Michigan Wastewater Dashboard for COVID-19 surveillance (MIwastewater) and to weekly percentages of Emergency Department (ED) visits diagnosed as COVID-19 in Michigan, as reported by the CDC COVID Data Tracker (CDCcovidtracker). Figure 3.8 demonstrates that our estimated trends between March 22, 2020, and October 24, 2022, closely align with wastewater-based surveillance for SARS-CoV-2 shed into Michigan’s public sewer systems. Both surveillance methods capture the spikes in November 2020 and January 2022, as well as the downward trend since August 2022. Notably, our surveillance approach can anticipate increases in ED visit numbers reported to the COVID Data Tracker by approximately one to two weeks. This validation supports findings previously reported in Indiana (mrp-covid21; mrp-covid22).
Our results based on a representation adjustment of routine hospital test records serve a synthetic proxy for random sampling. When available, random-sample testing surveys provide valuable benchmarking data and should be leveraged to calibrate other data sources to ensure population representativeness (irons2021estimating; menachemi2020population). However, increasing nonresponse rates in these surveys necessitate nonresponse bias adjustments (NRBA-cr2022; NRBA-long2022). For example, yiannoutsos2021bayesian have applied a similar method to MRP and adjusted for nonresponse bias in a randomized study of COVID-19 testing in Indiana, the response rate of which is 23.6%. Notably, the trends in calibrated new infection numbers reported by irons2021estimating are consistent with those seen in the MRP-adjusted hospital test incidence monitoring (mrp-covid21) between March 2020 and March 2021 in Indiana, particularly regarding the capture of infection spikes.
4 Discussion
With generalizability as the goal, the MRP method adjusts for selection bias and stabilizes small group estimates. We extend MRP to both cross-sectional and longitudinal data and facilitate subgroup trend analyses over time with a user-friendly interface. A key contribution of our MRP interface is its capacity to standardize statistical workflows and facilitate reproducibility. By integrating all analytic steps—from data preprocessing and model fitting to diagnostics and result visualization—into a comprehensive pipeline, the platform makes it straightforward for others to replicate findings and validate results on new datasets. The interface is built using open-source tools (Stan, R, Shiny) and is publicly available, supporting transparency and broad accessibility. By making MRP methods widely available without substantial coding requirements, the interface empowers a wider range of users to implement advanced statistical methods and strengthens training in reproducible population health analytics. These advantages position the interface as a foundational resource for standardized use and methodological rigor in future epidemic surveillance and related scientific studies.
The interface tracks the COVID-19 epidemic and delivers substantive findings in time, which is demonstrated using the Michigan health system data. The findings based on the MRP workflow have implications for public health monitoring and policy decision-making. By correcting sample selection bias and generating stable estimates for small demographic and geographic subgroups, the approach can identify populations at higher risk of infection or with lower access to healthcare resources. Health agencies can use subgroup-specific incidence estimates to allocate testing resources and direct targeted communication to communities with greater vulnerability. Tracking epidemiological trends using routine hospital data provides actionable guidance for predicting surges, anticipating clinical burden and responding swiftly to local outbreaks. Furthermore, the analytic pipelines support transparency and comparability in health reporting, which is vital for evidence-based policymaking and resource allocation.
Our approach has several key assumptions. First, we assume sample selection is ignorable, conditional on adjusted demographics and geography. The only factor determining the sample inclusion is the selection for elective surgical procedures. Elective surgery patients may differ from the broader community in unmeasured ways, such as healthcare access or overall health status. However, we need high-quality population data on healthcare use and health measures to adjust these potentially confounding factors. Second, we assume a constant ratio of asymptomatic to symptomatic SARS-CoV-2 infections within any demographic and geographic stratum. We use the estimated incidence based on asymptomatic test records to track the infection trend, but not the magnitude. The ratios may change values with new viral variants and immunity levels that are naturally-acquired or vaccine-induced. We have conducted sensitivity analyses by including the time indicators when the new viral was first detected in Michigan in the model and found that the estimated trends are similar, though the incidence rates have slightly shifted. We have also applied the MRP adjustment to viral IgG testing data of the same group of asymptomatic patients and validated the method using verified clinical metrics of viral and symptomatic disease incidence to show the expected biological correlation of these entities with the timing, rate, and magnitude of seroprevalence (mrp-covid22). Third, the model-based adjustment is subject to model misspecification. We use a Bayesian binomial model, and our subgroup estimates are robust across different outcome models. It is possible that the model fails to capture some data structure. We suggest users conduct thorough model diagnostics, such as the PPC and LOO-CV in our paper, and result validation. Additionally, the interface’s current focus is infection incidence estimation by subgroup, but it is extensible to other epidemiological parameters (e.g., effective reproduction number, infection fatality ratio), which would require integrating individual-level test data with aggregate case and mortality data using hierarchical Bayesian frameworks. Despite these limitations, the flexibility of the MRP approach allows its use with a variety of sampling methods and data sources, supporting broader applicability beyond the specific settings tested so far. Though post-epidemic testing is currently paused, the interface has broad applicability for other disease monitoring and data analyses with population representation.
With respect to our current sampling method for COVID-19 viral tracking, in accordance with accepted American Society of Anesthesiology standards, all preoperative patients in the hospital system are subjected to surgical risk evaluation. Hence, routine hospital testing is already implemented uniformly across the U.S., increasing the operational feasibility of our proposed surveillance system. Previous work used data from a community hospital in Indiana (mrp-covid21), and the results show differences between the states. As shown in our county-level estimates in Michigan, geographic variation can be substantial. Expanding to data from more states will further enhance national generalizability. Empowering users of varying backgrounds to analyze their own localized data effectively, the workflow can be directly adopted by local, state, and national health departments to ensure methodological consistency.
Originally developed in response to COVID-19, the MRP interface provides a foundation for broader epidemic surveillance and diverse applications in health and social science research. It accommodates time-varying and cross-sectional data, continuous and binary outcomes, and supports subgroup analyses across demographic and geographic domains. Users can specify models, priors, and poststratification data, and analyze probability sample surveys, non-probability samples, and multiple data sources. Future enhancements will address complex spatiotemporal structures (e.g., integrating real-time individual and aggregate testing data with autoregressive structures across transmission neighborhoods), custom prior distributions, and poststratification with incomplete population data. The MRP workflow is broadly applicable across a range of public health surveillance contexts. For example, recent measles and pertussis outbreaks in various parts of the U.S. underscore the need for flexible, reliable surveillance systems that can rapidly estimate disease prevalence and identify at-risk populations (cdc-measles; cdc-pertussis). By facilitating subgroup estimation and adjusting for selection bias, our MRP interface could be readily adopted for outbreak response and ongoing surveillance of many infectious diseases, where sample representativeness and rapid trend analysis are critical. The potential to use routinely collected hospital data, school outbreaks, or sentinel surveillance and combine them with population-level information makes our framework adaptable beyond COVID-19, reinforcing its generalizability for future epidemic preparedness and monitoring of emerging health threats.
In summary, the MRP interface advances statistical methods for epidemic surveillance by combining methodological rigor, computational accessibility, and practical impact. As public health agencies move toward more data-driven and adaptive response models, tools that enable granular, timely, and reproducible estimation will become increasingly essential. By establishing a foundation for standardized population health analytics, our approach aims to support ongoing efforts in epidemic preparedness, resource optimization, and evidence-based policy.
Acknowledgments
Data Availability
The patient records of the COVID-19 tests are confidential and require data use agreement in compliance with HIPAA privacy regulations for access. The interface includes synthetic example datasets for replication and demonstration purposes.
The appendix provides supplemental analysis results.
Appendix A Summary of geographic characteristics
@@
Figure A.1: Distributions of geographic characteristics based on the linked American Community Survey data in the catchment area.
Table A.1 gives the distributions of geographic characteristics.
Appendix B Model fitting results
The Stan fit summaries of Model A are displayed in Table B.1.